(1 point) Consider the linear system 3-1_3__3]; a. Find the eigenvalues and eigenvectors for the coefficient matrix. Vi = andI AM STRUGGLING HEAVILY ON THIS PROBLEM PLEASE HELP ME WITH THE CORRECT ANSWERS AND NEAT WORK

AND PLEASE ANSWER ALL ANSWER BOXES PLEASE PLEASE AND THANK YOU
Answer #1

2. 3 -S رو کے -3 Let A = 3 2 3 x 2 Then A-XI1=0 gives - 5 -3-2 or, (3-X) (- 3->) -(10) = 0 x² - 9+10 x2 = -1 eigen values areThe given system can be written as 3 2 5-3 NOW (31) 3ti (Cost tisint) (* 3 Cost - sint ti cost ti 3 sint - 5 cost - 1.5 sint